OpenStudy (princeevee):

Evaluate *(sqrt)*x+21 when x = 4, write the answer in simplified radical form ((will medal)) A. 25 B. 23 C. 10 D. 5

OpenStudy (princeevee):

\[\sqrt{21 + x} when x = 4 \]

OpenStudy (princeevee):

sorry if this dosent look right

OpenStudy (princeevee):

anyone? ;-;

OpenStudy (hysusonic):

dude what is 21+4 since x=4 and the equation is 21+x

OpenStudy (princeevee):

its \[\sqrt{21+x}\]

OpenStudy (princeevee):

X = 4

OpenStudy (hysusonic):

dude 21+4 is the same as 21+x because x=4 so you substitute x for 4

OpenStudy (princeevee):

can u help me ? : /

OpenStudy (hysusonic):

ugh it's 25 \[21+4=25\] simple addition

OpenStudy (princeevee):

isnt there supose to square root it?

OpenStudy (hysusonic):

that's after you substitute x for 4

OpenStudy (princeevee):

so 5?

OpenStudy (hysusonic):

YES PRAISE THE HIGH HEAVENS IF THEY EXIST YOU KNOW HOW TO ADD 21 AND 4

OpenStudy (princeevee):

.

OpenStudy (princeevee):

so is it 25 or 5 if you square root it

OpenStudy (anonymous):

for other reading this guy sucks why because he isn't patience.